12-13 October 2008

Observers: J. Hill, V. Knierim, W. Seifert, N. Ageorges and other LUCIFER folks, plus J. Schmidt working on DIMM
Software Support: M. Delapena, N. Cushing in Tucson
Telescope Operators: D.Gonzalez-Huerta, W.Wack

Summary

Clear weather at the beginning, but 2 arcsec seeing caused by the cold front that went through yesterday. We spent the evening discovering various obscure bugs as we exercised the system. We closed due to high winds after 11PM. After some HBS problems, we opened again at 4:15 AM.

Details

After several days of bad weather, the sky has cleared and we are able to open for sky flats. Ran SX mirror ventilation in the afternoon, so the mirror is not so far out as it might be. Butch added 30# of refrigerant to the summer chiller, so it is working better if not completely optimal.

Times reported in UT below.

Open for sky flats

00:25 Open for sky flats facing east at zenith. Sky is completely clear with a couple wisps of cirrus. Wind from West varying from 15-20 m/sec.

Preset to alignment star

01:30 Preset to the alignment star

01:45 Adjust IE=-168 CA=+2 (Telescope structure is at +7 degC, while ambient is at 0 degC.

That star was too bright even for 3 sec WFS exposures, so collimation wasn't converging well.

02:00 Preset to BS9183 gs=25. (failed because PA=0 not appropriate for this guide star)

02:07 Preset to BS9183 gs=25 PA=254
(02:08:05 preset command completed)

Collimation has converged OK, but note that Z4 is changing -150 nm/min. Seeing is 1.5-2 arcsec as seen on WFS apertures. WFS spots show the "usual" quadrifoil pattern when things are out of temperature equilibrium. SX primary glass temp is 3.7 degC.

02:30 Preset to adjust pointing again after collimation. IE=-195.83 CA=+16.82 Apparently the temperature gradient is still building in the telescope structure.

Problem with WFS exposure time

02:40 This Active preset had problems - it took a long time before taking any WFS image. Then it adjusted the WFS exposure time up to 120 sec and caused the WFS spots to saturate on this star. Apparently this strange exposure time happened earlier:

Oct 12 19:08:47 lbtmu107 LBT_left_GCS: WFS: peak value (381) above background (1099) is exceeding limits. Adjusting exposure time to 120000 ms.
This was the first WFS exposure of the preset that completed at 02:08:05 UT = wfsc 0350 = which had peak counts in the spots 240 counts above background. The second image with exposure increased to 120 sec = wfsc 0351 = had peak counts of 1000 counts above background. In this sense, GCS did what it was advertised to do. The actual problem came on the next preset when the exposure time was still 120 sec and we had a brighter star. Now in image wfsc 0358, the peak counts are 31000, but the spots are starting to bleed along columns. It seems that this CCD has a saturation (bleeding) level much lower than 65000 counts. See Issue #1718.

02:44 setWFSExpTime 30000 to manually force it back to nominal exposure. Then the collimation correction was OK.

Problem with GCS reading images from AzCam

02:53:20 Preset fails, seems like GCS never received the acquisition image.

02:59 Abort preset and stop GCSL

03:00:58 Preset again

03:02:51 readGuideCam -e 1000 hangs for more than 1 minute.

03:05 stop GCS, stopAGW, startAGW, start GCSL

20:06:22 readGuideCam -e 1000 Oct 12 20:06:22 lbtmu107 LBT_Networkserver: GCSL died unexpectedly

Now it works after restarting AzCamServers a second time, and restarting GCSL a third time. So we don't understand where the problem was. See Issue #1719.

Checking vignetting around the field

03:40 Nancy making presets to check the vignetting of the guide probe with respect to LUCIFER field and WFS vignetting.

03:54 Star on edge of acquisitiion field fails to appear on guider - this guide star was 4 arcmin off-axis

03:56 Preset in Parallactic mode to recheck the pointing. IE=-192.8 CA=12.0

04:01 Now the same star is at 310,160 on the acquisition image (better, but still not perfect) This time the guider pulls it in without problems.

04:19 With guide star 4.5 arcmin off-axis we have lost 1 row of spots at the edge of the pupil (bottom row from Doug's display, and top row from the GCS GUI display)

04:25 With guide star 5.0 arcmin off-axis we have lost 1.5 rows at the edge of the pupil, and also see 1 corner of the M3 vignetting - see wfsc 404.

04:30 Question for Doug: Nancy noticed on the IDL displays of WFS analysis that the illuminated spots cover rows 3-13 on wfscimage000412_dd.fits, but the arrows displayed are in rows 2-12. Is that a bug or a feature? In the X-direction illuminated spots cover columns 1-12, but arrows are displayed on columns 2-12.

Problem with spurious coma at the very edge of the field

04:41 With guide star 5.33 arcminutes off-axis, we have lost two full rows of WFS spots.

04:42 SX primary mirror has encountered the +X limit. Was this stimulated by a bad correction from these WFS images with 2 missing rows. It looks like we really ran into the static supports since image was distorted.

Andrew comments that he and Doug never encountered a problem like this during their off-axis collimation tests.

04:48 Clear Active Optics

04:49 Preset with guide star 4.5 arcmin off-axis. Now we are waiting for WFS to converge on the huge focus and coma corrections starting with wfsc 0419. Part of the reason that the convergence is slow is the fact that we have ~2.6 arcsec seeing as reported on GCSGUI. This sequence end with wfsc 0432.

05:00 SX Active Optics : X=-1.636 Y=-2.512 Z=-0.265 RX=-26.9 RY=17.6

05:04 Preset 05:20 arcmin off-axis. Starting with wfsc 0433 through wfsc 0436. Apparently the two missing rows are giving us a false signal of coma, which is not corrected by an actual coma adjustment of the mirror. Then with 3-5 application of this false coma signal, the mirror is moved -1.5 mm in X and runs into the limit.

05:12 Preset back on-axis. The thing I don't understand is that it sent Z8=+1500 to undo the effects of the false coma which appeared to be Z8=+280 applied a few times. ???

The problem seems to be that Torsten's code in GCS is getting a different answer than Doug's IDL code. The discrepancy seems to be for these particular vignetted images, and not for the Doug-Torsten comparison in general. Here are the results that each would have sent to M1 for image wfsc 0433.

Poly Doug Torsten
Z4 -397 -89
Z5 41 -50
Z6 -4 367
Z7 -77 -192
Z8 -173 -793

See Issue #1714.

Seeing continues to be in the range 2-2.5 arcsec as reported by the guider. Both primaries are within 1 degC of ambient, but the steel structure is warmer. Turn off mirror ventilation on DX to reduce load on chiller.

Problem with AGw Motor 1 encountering limit

05:50 Offset command hangs partway through. Probe is at X=175 Y=540, so it hasn't completed the move to Oct 12 22:50:51 lbtmu107 LBT_left_GCS: AGWUnitAIP::doSetProbePosition: calling oacontrol::setxy(10.144.0.64, 0, x:180.218435 y:540.173947, 0, 0) ...

After ~7 minutes, we aborted the offset from the IIFGUI, but Nancy still didn't get a return back through irs/LUCIFER.

I am not able to move the probe as agwuser:
setxy -x 180 -y 540
setxy: AGW1: Error moving the motor(s).
Aborting...

Close for Wind

06:05 Close for increasing wind. David measures 49.9 mph in the parking lot. At least this allows us to catch up on our Issuetraks.

06:46 Stop GCS, still can't move probe.

06:47 stopAGW, startAGW, but RPC communication error when homing motor.

06:49 stopAGW (errors in oacserver window)

06:51 kill oacserver and check4nports, and restart them. Then we are able to home the motors.

Kill IIF and restart. Michele had problems with this.

06:53 My manual command never returns: setxy -x 180 -y 540

07:02 kill and restart oacserver and check4nports again.

agwuser@agw-control:~$ setxy -x 100 -y 540
Command successfully transmitted!

agwuser@agw-control:~$ setxy -x 170 -y 540
Command successfully transmitted!

agwuser@agw-control:~$ setxy -x 175 -y 540
Command successfully transmitted!

agwuser@agw-control:~$ setxy -x 180 -y 540
hangs

Motor 1 is encountering a positive limit and causing the oacserver software to hang. See Issue #1717.

HBS Problems

10:00 Wind has dropped to ~15 m/sec, but now HBS is giving problems by shutting down.

David and Wayne reset everything related to HBS without much effect. My analysis is that there were multiple things wrong:

a) EL comp pressure E2 wasn't on the default force list in WinCC. We forced it.
b) Pump 1 seems to have flow problems (less flow than Pump 2) - We now have pump 1 turned off.
c) The default flow table is a bit too low at these ambient temperatures - slighly below 0 degC. We increased the flow by 10 units in each of table nodes 02 and 03.
d)The "stirring" pump was not running, and these are the temperatures where the hot return channel. Now the stirring pump is plugged in and running.

Open again

11:15 Open dome with HBS problems apparently solved.

11:18 Nancy's preset fails with invalid handle message. Restarted PCS and IIF, and PCSGUI.

11:24 Nancy's preset fails on the parkActiveProbe issue.

11:30 Preset succeeds in Track mode. But then the stepFocus command failed.

11:35 Preset succeeds. CA=15.0 IE=-180.

11:40 Preset in Active Mode, WFS adjusts exposure for 2 min images.

11:50 Off-axis Active Preset

11:52 setWFSExpTime 30000

Seeing is still in the 2-2.5 arcsec range.

11:55 Preset in track mode, to take images at different rotator angles.

SX Primary Active X=-1.732 Y=-2.672 Z=-0.420 RX=-28.7 RY=+18.6 EL=57 plus Zglobal=+0.8

12:19 Preset in Active Mode

Why did GCS wait 2 minutes between these two commands?
Oct 13 05:19:12 lbtmu107 LBT_left_GCS: WFS: requesting a WFS image...
Oct 13 05:21:12 lbtmu107 LBT_left_GCS: WFS: intending to wait 127716 ms for the WFS image.

It seems odd that it took 4.5 sec to NOT move the already parked probe!
Oct 13 05:33:08 lbtmu107 LBT_left_GCS: GCSL:parkActiveProbe executed
Oct 13 05:33:08 lbtmu107 LBT_left_GCS: didn't move probe as it was already in the correct position.
Oct 13 05:33:12 lbtmu107 LBT_left_GCS: GCSL:parkActiveProbe execution time: 4506 ms at 1223901192835
My suspicion is that lbtmu107 is having some swapping problems since the swap space in use decreased when I stopped my version of DDViewer. Also note that OSS has swelled to 315m of virtual memory, and GCSL to 350m.

Oct 13 05:42:44 lbtmu107 LBT_left_GCS: AGWUnit initialized, calling init() ...
Oct 13 05:42:46 lbtmu107 LBT_left_GCS: AGWunit init(): _filterPosition or _focus read back from AGW server is negative indicating an error.
Oct 13 05:42:46 lbtmu107 LBT_left_GCS: init() returned an error, indicating the AGW unit is not accessible.
Oct 13 05:42:46 lbtmu107 LBT_left_GCS: Couldn't change AGW to AIP_LUCIFER.
Oct 13 05:42:46 lbtmu107 LBT_left_GCS: GCSL:setGuidestarList execution time: 2422 ms at 1223901766670

agwuser@agw-control:~$ getfocus
Focus position: -0.001000
We moved the focus to +1 mm and everybody was fine. See Issue #1722.

However, by now things are getting too bright to do WFSing.

Morning sky flats

13:00 At zenith to take sky flats.

13:50 Close Dome. Sky is clear. Wind has dropped to 5 m/sec. T=-3degc, D=-15degC. -- JohnHill - 13 Oct 2008
Topic revision: r6 - 19 Oct 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback